# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v1

@@ 4.3 Massive Object Orbits and Precession
Table 1: The local minimum position of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM. In this table, the black hole electric charge and non-linear electrodynamics parameter are $Q=0.5M$, $a=0.5M^{2}$, and conserved angular momentum in the massive object’s bound orbit is selected as $J=8M$. In this table, we also substitute $1/2$ in the effective potential given in equation (9) such that the effective potential for massive object vanished at infinity $\lim_{r\to\infty}V_{\text{eff}}(r)=\lim_{r\to\infty}\big{\{}\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}-\frac{1}{2}\big{\}}=0$.
PFDM parameter $\lambda_{\text{DM}}=-0.5M$ $\lambda_{\text{DM}}=-0.4M$ $\lambda_{\text{DM}}=-0.2M$ $\lambda_{\text{DM}}=-0.1M$ $\lambda_{\text{DM}}=0.0M$
local minimum of $V_{\text{min}}=-0.03566$ $V_{\text{min}}=-0.03092$ $V_{\text{min}}=-0.01969$ $V_{\text{min}}=-0.01408$ $V_{\text{min}}=-0.00841$
effective potential
@@ 4.4 Black Hole Shadow
This subsection discusses the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM. The critical shadow radius of a black hole can be calculated from the effective potential of photons moving in gravitational field, or through a geometric approach proposed in a recent works [102]. Here, we calculate the radius of unstable photon sphere and the black hole shadow radius based on the geometric approach, in which the photon sphere and its stability are uniquely determined from Gaussian curvature and geodesic curvature in the optical geometry of black hole spacetime. Particularly, for any photon sphere, its geodesic curvature vanishes, which is completely equivalent to find the local extreme point of effective potential [102, 103, 104]
@@
The unstable photon spheres (or stable photon spheres) are determined by the negative (or positive) sign of Gaussian curvature, which is equivalent to find the local maximum (or local minimum) points of effective potential [102, 103, 104]
@@
Figure 6: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels highlight the influences coming from black hole electric charge, dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the critical shadow radius affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$ and $Q=0.5M$). (c) The middle left panel shows the critical shadow radius affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$). (d) The middle right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$ and $Q=0.5M$). (e) The lower left panel presents the critical shadow radius influenced by nonlinear electrodynamics effects (with $\lambda_{\text{DM}}=-0.2M$). (f) The lower right panel shows the black hole shadow in celestial coordinates influenced by black hole electric charge (with $a=0.5M^{2}$ and $\lambda_{\text{DM}}=-0.2M$). In all four panels of this figure, the black hole shadow radius, black hole electric charge, PFDM parameter and nonlinear electrodynamics parameter are given in unit of black hole mass $M$.
@@
The numerical results on black hole shadow for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 6. The left panels plot the critical shadow radius $r_{sh}$ calculated using expression (29), and the right panels illustrate the black hole shadow in celestial coordinates, highlighting the black hole shadow affected by three black hole parameters (the black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$). From figure 6, it is reasonable to draw the following conclusions. Firstly, the increasing of black hole electric charge diminishes the apparent size of black hole shadow. Secondly, the lower left panel shows that the nonlinear electrodynamics parameter $a$ has very tiny influences on the black hole shadow size. For the same electric charge and PFDM parameter, the critical shadow radius calculated with different nonlinear electrodynamics parameter values do not exhibit any obvious differences. Thirdly, when PFDM parameter is not vary large, such as $\lambda_{\text{DM}}\leq 0.5M$ in upper and middle panels of figure 6, the size of black hole shadow is enlarged as PFDM parameter reduced (but this statement is no longer valid when PFDM reaches a large critical point, see figure 7). Particularly, a negative PFDM parameter results in a much larger black hole shadow size.
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) do not produce any naked singularities (see Appendix A for the discussions on horizons of Euler-Heisenberg black hole surrounded by PFDM), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.2M,0.0M,-0.2M,-0.5M$, the critical shadow radius calculated using expression (29) is not continuous at a large black hole charge $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge excesses the critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change, and some branch of unstable photon spheres are disappeared. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is natural to question that the results obtained using expression (29) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
@@
To further explore the dark matter effects on black hole shadow, the figure 7 gives the critical shadow radius changed with respect to PFDM parameter, both in the presence and absence of nonlinear electrodynamics effects. It can be found that when PFDM parameter is smaller than a critical value ($\lambda_{\text{DM}}<\lambda_{\text{critical}}$), the black hole shadow radius reduced with the increasing of PFDM parameter. However, when PFDM parameter excesses this critical value ($\lambda_{\text{DM}}>\lambda_{\text{critical}}$), the black hole shadow radius could be magnified when PFDM parameter increases. This seems to be a universal properties of the PFDM effects on black hole shadows. Some recent studies on charged RN black hole surrounded by PFDM have reported the same conclusion on black hole shadow as we have presented here (the existence of critical point $\lambda_{\text{critical}}$) [50, 49, 51]. Furthermore, form figure 7, the critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values. Since the nonlinear electrodynamics effects have tiny influences on black hole shadow size, the left and right panels of figure 7 exhibit very similar results and tends.
@@ 5 Conclusions
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which are obtained from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics and PFDM effects in charge black hole system. It provide us a simply way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects influenced on charged black hole systems. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by black hole parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tend for various gravitational lensing observables (including the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits, and the black hole shadow size). Firstly, the Euler-Heisenberg black hole with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these gravitational lensing observables. Thirdly, the most significant impacts on these gravitational lensing observables coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, the PFDM with a negative dark matter parameter could greatly enhance the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM plays a role of “effective mass”. A negative PFDM parameter contributes to a positive “effective mass” in the spacetime metric, therefore it can enhance the gravitational lensing observables (such as gravitational deflection angle, time delay of light, and precession angle of massive object’s orbits). Moreover, the dark matter medium also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow, when PFDM parameter is smaller than a critical value, the black hole shadow radius reduced as the increasing of PFDM parameter. When PFDM parameter excesses this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
@@
The results in our present work can deepen our understanding of the nonlinear electrodynamics effects / QED effects and dark matter effects on charged black hole systems in the gravitational lensing. Hopefully it may have potential applications for theoretical and observational probing for these dark matter effects and nonlinear electrodynamics effects in black holes and gravitation. Since the dark matter could significantly influence the gravitational lensing observables, the PFDM parameter can be constrained from gravitational lensing observations, by analyzing and comparing the observed data (on gravitational deflection angles, time delays of light, precession angle of bound orbits, apparent black hole shadows) with theoretical predictions, especially with the help of accumulating observational data for supermassive black hole in galaxy centers. However, the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables, so it would be not easy to probe the nonlinear electrodynamics in gravitational lensing observations (unless very high-precision observational data are employed).
@@ Acknowledgment
The authors thanks Song-Lin Lyu and Guang-Zhou Guo for helpful discussion and comments on the numerical schemes. This research was funded by the Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126), the Scientific Research Program of Chongqing Science and Technology Commission (the Chongqing “zhitongche” program for doctors, Grant No. CSTB2022BSXM-JCX0100), the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2020ZDZ027) and the Research and Innovation Team Cultivation Program of Chongqing University of Technology (Grant No. 2023TDZ007).
Abbreviations
The following abbreviations are used in this manuscript:
QED Quantum Electrodynamics
PFDM Perfect Fluid Dark Matter
RN Reissner-Nordström
WIMPs Weakly Interactive Massive Particles
EHT Event Horizon Telescope
@@
Appendix A Horizon of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
@@
This appendix summarizes the results on event horizon of Euler-Heisenberg black hole surrounded by PFDM. The positions of horizon can be solved from the metric component function
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}=0$ (30)
@@
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a=0$ and $\lambda_{\text{DM}}$), the spacetime reduce to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects (or QED effects) and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which are calculated by numerically solving equation (30).
@@
Figure 8: The horizons of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels highlight the influences coming from dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the charged black hole’s horizons affected by PFDM parameter $\lambda_{\text{DM}}$ without the nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel shows the Euler-Heisenberg black hole’s horizons affected by nonlinear electrodynamics effects (parameterized by $a$) without the PFDM. (c) The lower left panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.2M^{2}$). (d) The lower right panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.8M^{2}$). In all four panels of this figure, the horizon radius, black hole electric charge, PFDM parameter and nonlinear electrodynamics parameter are given in unit of black hole mass $M$.
@@
From the upper left panel of figure 8, it can be observed that the PFDM does not affects the structure of horizons in the absence of nonlinear electrodynamics effects (with $a=0$). Without the nonlinear electrodynamics effects / QED effects, the charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. The extremal black hole charge $Q_{\text{extreme}}$ become increased for a smaller PFDM parameter $\lambda_{\text{DM}}$. This can be explained from the “effective mass” role that PFDM played in the spacetime metric (6), as we have explained in section 4. A negative PFDM parameter ($\lambda_{\text{DM}}<0$) could contribute to a positive “effective mass” in the spacetime metric, therefore the corresponding extremal black hole charge can be greater than black hole mass $M$ for a negative PFDM parameter. When the black hole charge excesses the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizon-less and produces a naked singularity in the center point $r=0$.
@@
For Euler-Heisenberg black holes in the presence nonlinear electrodynamics effects (with $a>0$), the structure of horizons can be changed, which are clearly observed in the upper right panel of figure 8. Firstly, the naked singularity no longer exists for large black hole charge when nonlinear electrodynamics is considered. For any given black hole charge, there are at least one horizons near Euler-Heisenberg black hole. Secondly, for a small black hole charge, the spacetime has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel, various curves of horizon radius correspond to different nonlinear electrodynamics parameter values $a$ in the branch $r_{H}\sim 2M$ and $Q<M$ are almost overlapped with each other in the upper right panel, but only the pure RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black dotted curve). Thirdly, for black hole with a relative intermediate electric charge ($0.85M<Q<1.05M$), the Euler-Heisenberg black hole could have three horizons. However, it should be mentioned that the Euler-Heisenberg black hole with a very large nonlinear electrodynamics parameter $a$ (for instance the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons, producing a single horizon for all electric charge values.

### v2

@@ 4.3 Massive Object’s Bound Orbits and Precession Angle
Table 1: The local minimum position of effective potential for massive objects moving around Euler-Heisenberg black hole in the presence of PFDM. In this table, the black hole electric charge and nonlinear electrodynamics parameter are chosen to be $Q=0.5M$, $a=0.5M^{2}$, and conserved angular momentum for massive object’s bound orbit is selected as $J=4M$. In this table, we have substituted $1/2$ in the effective potential given in equation (10) such that the effective potential for massive object vanished in the asymptotic flat region (the infinity) $\lim_{r\to\infty}V_{\text{eff}}(r)=\lim_{r\to\infty}\big{\{}\frac{f(r)}{2}\big{[}\frac{J^{2}}{r^{2}}+\epsilon\big{]}-\frac{1}{2}\big{\}}=0$.
PFDM parameter $\lambda_{\text{DM}}=0.0M$ $\lambda_{\text{DM}}=0.05M$ $\lambda_{\text{DM}}=0.1M$ $\lambda_{\text{DM}}=0.2M$
local minimum of $V_{\text{eff}}$ $V_{\text{min}}=-0.03611$ $V_{\text{min}}=-0.02539$ $V_{\text{min}}=-0.01807$ $V_{\text{min}}=-0.00821$
@@ 4.4 Black Hole Shadow
This subsection discusses the black hole shadow for Euler-Heisenberg black hole surrounded by PFDM. The critical shadow radius of a black hole is determined by the unstable photon sphere, which can be calculated from the effective potential of photons moving in gravitational field, or through a geometric approach proposed in a recent works [111]. Here, we calculate the radius of unstable photon sphere and the black hole shadow radius based on the geometric approach, in which the photon sphere and its stability are uniquely determined from Gaussian curvature and geodesic curvature in the optical geometry of black hole spacetime. Firstly, any photon orbits maintain their geodesic nature when they are transformed into optical geometry, so the geodesic curvature for a photon sphere vanishes. This geodesic curvature condition for photon sphere is completely equivalent to find the local extreme point of effective potential [111, 112, 113]
@@
Furthermore, the unstable photon spheres (or stable photon spheres) are determined by the negative (or positive) sign of Gaussian curvature, which is equivalent to find the local maximum (or local minimum) points of effective potential [111, 112, 113]
@@
Figure 6: The black hole shadow of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels of this figure highlight the influences coming from black hole electric charge, dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the critical shadow radius affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the absence of nonlinear electrodynamics effects (with $a=0$ and $Q=0.5M$). (c) The middle left panel shows the critical shadow radius affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$). (d) The middle right panel illustrates the black hole shadow in celestial coordinates affected by PFDM in the presence of nonlinear electrodynamics effects (with $a=0.5M^{2}$ and $Q=0.5M$). (e) The lower left panel presents the critical shadow radius influenced by nonlinear electrodynamics effects (with $\lambda_{\text{DM}}=0.2M$). (f) The lower right panel shows the black hole shadow in celestial coordinates influenced by black hole electric charge (with $a=0.5M^{2}$ and $\lambda_{\text{DM}}=0.2M$). In all panels of this figure, the black hole shadow radius $r_{sh}$, black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$ are given in unit of black hole mass $M$.
@@
The numerical results on black hole shadow for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 6. The left panels plot the critical shadow radius $r_{sh}$ calculated using expression (30), and the right panels illustrate the black hole shadow in celestial coordinates, highlighting the black hole shadow affected by three black hole parameters (the black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$). From figure 6, it is reasonable to draw the following conclusions. Firstly, the increasing of black hole electric charge diminishes the apparent size of black hole shadow. Secondly, the lower left panel shows that the nonlinear electrodynamics parameter $a$ has very tiny influences on the black hole shadow size. For the same electric charge and PFDM parameter, the critical shadow radius calculated with different nonlinear electrodynamics parameter values do not exhibit any obvious differences. Thirdly, when PFDM parameter is not vary large, such as $\lambda_{\text{DM}}\leq 0.5M$ in upper and middle panels of figure 6, the size of black hole shadow is enlarged as PFDM parameter reduced. However, this statement is no longer valid when PFDM reaches a large critical point, see figure 7.
@@
To further explore the dark matter effects on black hole shadow size, we use figure 7 to show the critical shadow radius changed with respect to PFDM parameter, both in the presence and absence of nonlinear electrodynamics effects. It can be observed that when PFDM parameter is smaller than a critical value ($\lambda_{\text{DM}}<\lambda_{\text{critical}}$), the black hole shadow radius reduces with the increasing of PFDM parameter. However, when PFDM parameter excesses this critical value ($\lambda_{\text{DM}}>\lambda_{\text{critical}}$), the black hole shadow radius get magnified as the increasing of PFDM parameter. This seems to be a universal properties of the PFDM effects on black hole shadows. Recent studies on charged RN black hole surrounded by PFDM have reported similar conclusions as we have presented here (the existence of critical point $\lambda_{\text{critical}}$) [50, 49, 51]. Furthermore, figure 7 shows that the critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values. Since the nonlinear electrodynamics effects have tiny influences on black hole shadow size, the left and right panels of figure 7 exhibit very similar results and tends.
Interestingly, the black hole shadow results for Euler-Heisenberg black hole surrounded by PFDM in the presence and absence of nonlinear electrodynamics effects / QED effects could also exhibit a slight difference. In the absence of nonlinear electrodynamics effects, because of the emergence of naked singularity when black hole charge $Q>Q_{\text{extreme}}$, the black hole shadow disappears for large black hole electric charge. On the other hand, in the presence of nonlinear electrodynamics effects, although the Euler-Heisenberg black hole (surrounded by PFDM) do not produce any naked singularities (see Appendix A for the discussions on horizons), the black hole shadow may still disappear under some circumstances. In the middle left panel of figure 6, when PFDM parameter takes the values $\lambda_{\text{DM}}=0.0M,0.1M,0.2M$, the critical shadow radius calculated using expression (30) is not continuous when black hole charge reaches a critical value $Q=Q_{\text{discontinuity}}$. This is caused by the discontinuity of unstable photon sphere positions in the vicinity of Euler-Heisenberg black hole. When black hole charge excesses this critical value, the number of unstable photon sphere around Euler-Heisenberg black hole undergoes a sudden change  [^3] When black hole charge satisfies $Q<Q_{\text{discontinuity}}$, there are two unstable photon spheres near the Euler-Heisenberg black hole (surrounded by PFDM), and the outer branch of unstable photon sphere determines the black hole shadow size via expression (30). When black hole charge becomes $Q>Q_{\text{discontinuity}}$, there is only one unstable photon sphere. The outer branch of unstable photon spheres in the $Q<Q_{\text{discontinuity}}$ cases is disappeared, which makes the unstable photon sphere position undergo a sudden change when black hole charge come across the critical point $Q=Q_{\text{discontinuity}}$.. Once the black hole charge comes across the discontinuity point (namely $Q>Q_{\text{discontinuity}}$), it is questionable that the results obtained using expression (30) could still represent a conventional “black hole shadow” or not. This issue deserves a more deeply study in the future.
@@ 5 Conclusions
In this work, the gravitational lensing of Euler-Heisenberg black hole surrounded by PFDM is studied. This kind of black hole solution, which are solved from the gravitational field equation coupled with nonlinear electromagnetic field and dark matter field, could give rise to the nonlinear electrodynamics effects and PFDM effects in charge black hole systems. It provide us a simply way to investigate the interplay of nonlinear electrodynamics effects (or QED effects) and dark matter effects on the studies of black holes. These effects coming from nonlinear electrodynamics and dark matter can be parameterized by parameters $a$ and $\lambda_{\text{DM}}$ respectively. In the presented work, we mainly focus on the gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angle of light, time delay of light, precession angle and trajectories of massive object’s bound orbits, black hole shadow are calculated numerically from geodesics.
@@
Our numerical results presented in this work show that influences coming from black hole electric charge, nonlinear electromagnetic effects, dark matter effects exhibit a similar tend for various gravitational lensing observables (including the gravitational deflection angle of light, time delay of light, precession angle of massive object’s bound orbits, and the black hole shadow size). Firstly, the Euler-Heisenberg black hole surround by PFDM with a larger electric charge could reduce the gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius. Secondly, the nonlinear electrodynamics effects parameterized by $a$ has very tiny influences on these gravitational lensing observables. Thirdly, the most significant impacts on these gravitational lensing observables coming from the dark matter effects. The gravitational deflection angle, time delay of light, precession angle of bound orbits and black hole shadow radius are all greatly changed when PFDM parameter $\lambda_{\text{DM}}$ is varied. Particularly, a larger PFDM parameter could greatly reduce the gravitational deflection angle, time delay of light and precession angle of bound orbits. This can be explained from the spacetime metric function (6) for Euler-Heisenberg black hole surrounded by PFDM, in which the PFDM parameter plays a role of “effective mass”. A positive PFDM parameter contributes to a negative “effective mass” in the spacetime metric, therefore it can diminish the gravitational lensing observables. Moreover, the dark matter also has notable influences on the shape and eccentricity of massive object’s bound orbits. For black hole shadow size, when PFDM parameter is smaller than a critical value, the black hole shadow radius reduced as the increasing of PFDM parameter. When PFDM parameter exceeds this critical value, the black hole shadow radius could be magnified when PFDM parameter increases. The critical value of PFDM parameter is roughly $0.5M<\lambda_{\text{critical}}<0.8M$ for different black hole charge values.
@@
We hope that the results in our present work can deepen our understanding of the nonlinear electrodynamics effects / QED effects and dark matter effects on the gravitational lensing of charged black hole systems. Hopefully it may have potential applications for theoretical and observational probing for dark matter effects in black holes and gravitation. Since the dark matter could significantly influence the gravitational lensing observables, the PFDM parameter can be constrained from gravitational lensing observations, by analyzing and comparing the observed data with theoretical predictions, especially with the help of accumulating observational data for supermassive black hole in galaxy centers. However, the nonlinear electrodynamics effects have very tiny influences on these gravitational lensing observables, so it would be not easy to probe the nonlinear electrodynamics in gravitational lensing observations (unless very high-precision observational data are employed).
@@ Acknowledgment
The authors thanks Song-Lin Lyu and Guang-Zhou Guo for helpful discussion and comments on the numerical schemes. This research was funded by the Natural Science Foundation of Chongqing Municipality (Grant No. CSTB2022NSCQ-MSX0932), the Scientific and Technological Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202201126 and No. KJQN202301164), the Scientific Research Program of Chongqing Science and Technology Commission (the Chongqing “zhitongche” program for doctors, Grant No. CSTB2022BSXM-JCX0100), the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2020ZDZ027) and the Research and Innovation Team Cultivation Program of Chongqing University of Technology (Grant No. 2023TDZ007).
@@
Appendix A Horizons for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
@@
This appendix presents results on horizons for Euler-Heisenberg black hole surrounded by PFDM. The radii of horizons can be solved from the metric component function
$f(r_{H})=1-\frac{2M}{r_{H}}+\frac{Q^{2}}{r_{H}^{2}}-\frac{aQ^{4}}{20r_{H}^{6}}+\frac{\lambda_{\text{DM}}}{r_{H}}\cdot\ln\frac{r_{H}}{|\lambda_{\text{DM}}|}=0$ (31)
@@
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a=0$ and $\lambda_{\text{DM}}=0$), the spacetime reduce to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which are calculated by numerically solving equation (31).
@@
Figure 8: The horizons of Euler-Heisenberg black hole surrounded by PFDM. The multiple panels highlight the influences coming from dark matter effects and nonlinear electrodynamics effects. (a) The upper left panel shows the charged black hole’s horizons affected by PFDM parameter $\lambda_{\text{DM}}$ without the nonlinear electrodynamics effects (with $a=0$). (b) The upper right panel shows the Euler-Heisenberg black hole’s horizons affected by nonlinear electrodynamics effects (parameterized by $a$) without PFDM. (c) The lower left panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.2M^{2}$). (d) The lower right panel plots the horizon radius of Euler-Heisenberg black hole in the presence of dark matter and nonlinear electrodynamics (with $a=0.8M^{2}$). In all four panels of this figure, the horizon radius, black hole electric charge $Q$, PFDM parameter $\lambda_{\text{DM}}$ and nonlinear electrodynamics parameter $a$ are given in unit of black hole mass $M$.
@@
From the upper left panel of figure 8, in the absence of nonlinear electrodynamics effects (with $a=0$), it can be observed that dark matter does not have obvious influences on the structure of horizons. Without the nonlinear electrodynamics effects, charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. When the black hole charge excesses the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizon-less and produces a naked singularity in the center $r=0$. In the presence of PFDM, the extremal black hole charge $Q_{\text{extreme}}$ become decreased than the conventional RN case. This can be explained from the “effective mass” role that PFDM played in spacetime metric (6), as we have explained in section 4. A positive PFDM parameter contributes to a negative “effective mass” in spacetime metric, thus the corresponding extremal black hole charge is less than black hole mass ($Q_{\text{extreme}}<M$) for a given positive PFDM parameter ($\lambda_{\text{DM}}>0$).
@@
In the presence nonlinear electrodynamics effects (with $a>0$), the horizon structure for charged Euler-Heisenberg black holes can be dramatically changed, which are clearly shown in the upper right panel of figure 8. Firstly, the naked singularity no longer exists when nonlinear electrodynamics is included ($a>0$). For any given black hole charge, there is at least one horizon in Euler-Heisenberg black hole. Secondly, for a small black hole charge, the Euler-Heisenberg black hole has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel of figure 8, various curves correspond to positive nonlinear electrodynamics parameter values ($a>0$) in the branch $M<r_{H}<2M$ and $Q<M$ are almost overlapped with each other, but only the pure RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black solid curve in the upper right panel). Thirdly, if the black hole electric charge fall into region $0.85M<Q<1.02M$, the Euler-Heisenberg black hole with a small nonlinear electrodynamics parameter ($a<0.5M^{2}$) could have three horizons. However, a larger nonlinear electrodynamics parameter (for instance the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons in Euler-Heisenberg black hole, producing a single horizon for arbitrary electric charge values.